OpenStudy (anonymous):

solve: (b+4)/(b+5)=(b+9)/(b+6)

OpenStudy (a_clan):

use componendo and dividendo

OpenStudy (a_clan):

by using c and d, you will simply get a linear equation instead of quadratic

OpenStudy (anonymous):

cross multiply b^2+10x+24=b^2+14x+45 -4x=21 x=-5.25